Kaity built a slingshot and put stone in it. It fires off the Science Center (45 m high) with a horizontal velocity of 20 m/s. When the stone hits the ground, how far from the building is the stone?
1
Expert's answer
2021-09-17T09:22:45-0400

Let's first find the time that stone takes to reach the ground:


"h=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2h}{g}}=\sqrt{\dfrac{2\cdot45\ m}{9.8\ \dfrac{m}{s^2}}}=3.03\ s."

Finally, we can find the distance from the building to the point where the stone lands:


"x=v_{0x}t=20\ \dfrac{m}{s}\cdot3.03\ s=60.6\ m."